Ms. A invests $400,000, $60,000, and $25,000 in stocks X, Y, and Z, respectively. The betas of these stocks are.25, .95, and 1.63, respectively. The market's expected return is 8%, and the risk-free rate is 2%. What is the expected value of Ms. A's investment?
Answer #1

Answer:

Total Investment = Investment in Stock X + Investment in Stock Y + Investment in Stock Z
Total Investment = $400,000 + $60,000 + $25,000
Total Investment = $485,000

Weight of Stock X = Investment in Stock X / Total Investment
Weight of Stock X = $400,000 / $485,000
Weight of Stock X = 0.82474

Weight of Stock Y = Investment in Stock Y / Total Investment
Weight of Stock Y = $60,000 / $485,000
Weight of Stock Y = 0.12371

Weight of Stock Z = Investment in Stock Z / Total Investment
Weight of Stock Z = $25,000 / $485,000
Weight of Stock Z = 0.05155

Portfolio Beta = Weight of Stock X * Beta of Stock X + Weight of Stock Y * Beta of Stock Y + Weight of Stock Z * Beta of Stock Z
Portfolio Beta = 0.82474 * 0.25 + 0.12371 * 0.95 + 0.05155 * 1.63
Portfolio Beta = 0.408

Portfolio Expected Return = Risk-free Rate + Portfolio Beta * (Market Return - Risk-free Rate)
Portfolio Expected Return = 0.02 + 0.408 * (0.08 – 0.02)
Portfolio Expected Return = 0.02 + 0.0245
Portfolio Expected Return = 0.0445